OpenStudy (anonymous):

evaluate the expression 32^-2/5

OpenStudy (anonymous):

\(32^{-\frac{2}{5}}\) The numerator in a fractional exponent is the power you raise the base to. The denominator in a fractional exponent is the root. The negative exponent means to invert the base.

OpenStudy (anonymous):

(32)^-2/5 = (2^5)^-2/5 = 2^-2 = 1/2^2 =1/4

OpenStudy (goformit100):

"Welcome to OpenStudy. I can answer your questions or guide you. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."